Solve the equations by the use of the operator D







D²y -6Dy +9y= e^3x + e^-3x



1
Expert's answer
2021-07-27T15:04:02-0400

"D^2y-6Dy+9y=e^{3x}+e^{-3x}"

For yn

"(D^2-6D+9)y=0\\m^2-6m+9=0"


"\implies(m-3)^2=0\\\implies m=3,3"


"\\yn=\frac{1}{D^2-6D+9}(e^{3x}+e^{-3x})"


"\\=\frac{e^{3x}}{D^2-6D+9}+\frac{e^{-3x}}{D^2-6D+9}"


"=\frac{xe^{3x}}{2D-6}+\frac{e^{-3x}}{36}"


So solution is "y(x)=y_n+y_p"


"\implies y(x)=ae^{3x}+c_2xe^{3x}+\frac{x^2e^{3x}}{2}+\frac{e^{-3x}}{36}"